### v1

# Cosmological perturbations of a relativistic MOND theory

Abstract
A relativistic MOND theory, promising in reproducing cosmology as well as the MOND phenomenology in the low acceleration regime, was recently proposed. We present the post-Newtonian (PN) approximation and relativistic perturbation equations of this theory in cosmological context. The PN equations are presented to 1PN order and perturbation equations are presented in fully-nonlinear and exact forms. The gauge issues are clarified. The 1PN equations and linear perturbation equations are presented without imposing temporal gauge conditions. The MOND field can be interpreted as a fluid with specific equation of state without anisotropic stress, and the Jeans criterion is derived for the MOND field.
I Introduction
Milgrom, in 1983, proposed MOND (Modified Newtonian Dynamics) to explain the gravity anomaly in galaxy rotations MOND-1983a . Instead of a popular way of saving conventional gravity (both Newton’s and Einstein’s) by introducing unseen dark matter (DM), he has modified in a non-relativistic context either gravity or inertia for the anomaly. Despite the lack of agreed relativistic theory, thus unable to address the cosmological demands other than the galactic rotation and related phenomena on similar scales, the non-relativistic proposal has survived till today as an alternative to the dark matter hypothesis. Recently, it gained independent observational support and challenge in stellar and Solar system scale phenomena; see Chae-2024 ; Hernandez-2024 for support from the long period binary stars observed by GAIA, and Vokrouhlicky-2024 ; Desmond-2024 for against in explaining dynamics and gravitational field in the solar system.
Such independent tests of the proposal in stellar and solar system scales are possible because Milgrom has proposed modification of gravity or inertia in the low acceleration regime by $f(|\vec{a}|/a_{M})\vec{a}=\vec{g}_{N}$ where $\vec{g}_{N}$ is the Newtonian gravitational acceleration, with $f(x)=1$ for $x\gg 1$, thus recovering Newton’s gravity, and $f(x)=x$ for $x\ll 1$ in the low acceleration MOND regime. In the MOND regime, we have $|\vec{a}|=\sqrt{a_{M}g_{N}}$ with $g_{N}=|\vec{g}_{N}|$. For an isolated spherically symmetric system with $g_{N}=GM/r^{2}$, using $|\vec{a}|=v^{2}/r$, we have $v=(GMa_{M})^{1/4}\propto M^{1/4}$ thus naturally recovering the observationally known galaxies’ flat rotation curves and Tully-Fisher/Faber-Jackson relation Sanders-McGaugh-2002 .
Observations have pin down the Milgrom’s universal constant of nature $a_{M}\simeq 1.2\times 10^{-8}cm/sec^{2}$ McGaugh-2004 . As originally noticed by Milgrom MOND-1983a the transition acceleration $a_{M}$ is intriguingly close to the acceleration scale of the expanding Universe $cH_{0}/(2\pi)=1.1\times 10^{-8}cm/sec^{2}$ using the current value of Lemaître-Hubble parameter $H_{0}=70(km/sec)/Mpc$. The interpolating function $f$ is not determined yet.
Bekenstein and Milgrom, in 1984, proposed a non-relativistic Lagrangian formulation of MOND, often called as AQUAL, leading to a modified Poisson’s equation $\nabla\cdot(f\nabla\Phi)=4\pi G\varrho$ with $\vec{a}=\nabla\Phi$ Bekenstein-Milgrom-1984 . At large distance from an isolated body with high symmetry, Milgrom’s original proposal is recovered Bekenstein-Milgrom-1984 . In the MOND based on the modified Poisson’s equation, the internal gravitational dynamics of a system is affected by an external gravitational field, often called the external-field effect, thus violating the strong equivalence principle Bekenstein-Milgrom-1984 . Variety of modifications of theory producing the MOND phenomenology is suggested Milgrom-2010-QUMOND ; Milgrom-2023-TriMOND ; Milgrom-2023-GQUMOND , and diverse relativistic extensions are also made Bekenstein-Milgrom-1984 ; Famaey-2012 .
Blanchet and Marsat Blanchet-Marsat-2011 introduced a small addition of a field in Einstein’s gravity to recover the MOND phenomenology in weak gravity low acceleration regime. They used a function of the acceleration of a four-vector $U_{a}$ associated with a scalar function $\tau$ as $U_{a}\propto\nabla_{a}\tau$. The theory is motivated by a quantum gravity proposal with local Lorentz invariance violation Horava-2009 ; Blas-2010 and Einstein aether theories Jacobson-Mattingly-2001 ; Jacobson-2010 ; Sanders-2011 .
The MOND proposal is successful in explaining the phenomenology in the galactic scale without demanding DM. However, successful cosmology also demands the presence of a sort of DM (a fluid with dust-like equation of state) in explaining phenomena in the large-scale structure, the cosmic background radiation (CMB), and the background matter content. Blanchet and Skordis Blanchet-Skordis-2024 recently proposed one more additional function in Blanchet and Marsat’s model, now using a function of the normalization amplitude of the same four-vector $U_{a}$ to accommodate such demands in cosmology.
The theory introduces two free functions of the $\tau$-field (MOND field): a function of the acceleration ${\cal J}(A)$ of a unit four-vector $U_{a}$ for the MOND phenomenology in low acceleration regime and the other a function of the normalization amplitude ${\cal K}(Q)$ of the same four-vector for successful cosmology, see Eq. (1). Forms of the two functions are still free to accommodate the phenomena in cosmology and MOND Blanchet-Skordis-2024 . Future studies may guide or design the functional forms of ${\cal J}(A)$ and ${\cal K}(Q)$ using theories and observations. For such studies the cosmological background and perturbation equations are needed.
Here, we present equations for the PN approximation and relativistic perturbation theory in cosmological context. The PN approximation is presented to 1PN order, and the perturbation theory is presented in fully-nonlinear and exact form. We address the gauge issues in both formulations and the 1PN equations and linear perturbation equations are presented without imposing temporal gauge conditions. These subjects are already studied in Blanchet-Skordis-2024 and here we elaborate some aspects. For both formulations, we compare the MOND contribution with an ordinary fluid. The MOND field does not contribute to the vector- and tensor-type perturbations, and it can be interpreted as a fluid with a specific equation of state (depending on the functional forms of ${\cal J}$ and ${\cal K}$) and is not associated with the anisotropic stress.
Section II summarizes the Lagrangian formulation of the theory with Einstein’s equation and the equation of motion in covariant forms. Sections III-V derive cosmological background equations, the cosmological PN equations to 1PN order, and the cosmological perturbation equations valid in fully-nonlinear and exact form, respectively. Section VI is a discussion.
In the post-Newtonian approximation we use $x^{0}=ct$ for derivation, whereas in the background and perturbation theory we use the conformal time $x^{0}=\eta$ with $cdt=ad\eta$; $a$ is the cosmic scale factor.
II Relativistic MOND
Blanchet, Marsat and Skordis Blanchet-Marsat-2011 ; Blanchet-Skordis-2024 introduced the following Lagrangian as a relativistic extension of the MOND proposal using a $\tau$-field
${\cal L}=\sqrt{-g}\Big{\{}{c^{4}\over 16\pi G}\big{[}R-2{\cal J}(A)+2{\cal K}(Q)\big{]}+L_{\rm m}\Big{\}},$ (1)
where
$\hskip-22.76228ptU_{a}\equiv-{c\over Q}\nabla_{a}\tau,\quad Q\equiv c\sqrt{-(\nabla^{a}\tau)\nabla_{a}\tau},$
$\hskip-22.76228ptA_{a}\equiv c^{2}U_{a;b}U^{b}=-c^{2}q^{b}_{a}\nabla_{b}\ln{Q},\quad A\equiv{1\over c^{4}}A^{c}A_{c},$ (2)
and $L_{m}$ is the ordinary matter (baryon and photon) part Lagrangian. Although we call the fundamental field simply a $\tau$-field, what is introduced is a four-vector field $U_{a}$ defined using the $\tau$-field; $U_{a}$ is a normalized timelike four-vector with $U^{c}U_{c}\equiv-1$; $q_{ab}\equiv g_{ab}+U_{a}U_{b}$ is a spatial projection tensor orthogonal to $U_{a}$; $A_{a}$ is an acceleration of the four-vector. For $\tau=\tau(t)$, we have $U_{a}=n_{a}$ where $n_{a}$ is the normal four-vector with $n_{i}\equiv 0$.
The ${\cal J}$-term is used to realize the MOND phenomenology in low acceleration regime by Blanchet and Marsat Blanchet-Marsat-2011 . The ${\cal K}$-term is added in order to accommodate the cosmological demands of the dark matter in CMB and the large-scale structure using the same $\tau$-field by Blanchet and Skordis Blanchet-Skordis-2024 . We may call it Blanchet-Marsat-Skordis (BMS) theory. The cosmological constant can be absorbed to ${\cal J}$ or ${\cal K}$ as ${\cal J}=\Lambda$ or ${\cal K}=-\Lambda$; here we consider the cosmological constant separately in $L_{m}$.
Variation with respect to $g_{ab}$ gives
$G_{ab}={8\pi G\over c^{4}}T_{ab},$ (3)
where the ${\cal J}$ and ${\cal K}$-parts are
$\hskip-14.22636ptT^{\cal J}_{ab}={c^{4}\over 8\pi G}\Big{[}-{\cal J}g_{ab}+{2\over c^{4}}{\cal J}_{,A}A_{a}A_{b}-{2\over c^{2}}({\cal J}_{,A}A^{c})_{;c}U_{a}U_{b}\Big{]},$
$\hskip-14.22636ptT^{\cal K}_{ab}={c^{4}\over 8\pi G}({\cal K}g_{ab}+Q{\cal K}_{,Q}U_{a}U_{b}).$ (4)
We will call the sum of these two parts the $\tau$-field (or $\tau$-fluid) energy-momentum tensor. We have $T_{ab}=T_{ab}^{m}+T_{ab}^{\tau}$ with $T_{ab}^{m}$ the matter part. The covariant fluid quantities measured by an observer with a general four-vector $u_{a}$ are introduced as Ellis-1971
$\hskip-28.45274ptT_{ab}=\mu u_{a}u_{b}+p\left(g_{ab}+u_{a}u_{b}\right)+q_{a}u_{b}+q_{b}u_{a}+\pi_{ab},$ (5)
$\hskip-28.45274pt\mu=T_{ab}u^{a}u^{b},\quad p={1\over 3}T_{ab}h^{ab},\quad q_{a}=-T_{cd}u^{c}h^{d}_{a},$
$\hskip-28.45274pt\pi_{ab}=T_{cd}h^{c}_{a}h^{d}_{b}-ph_{ab},$ (6)
where $h_{ab}\equiv g_{ab}+u_{a}u_{b}$ is a spatial projection tensor orthogonal to $u_{a}$.
Variation with respect to $\tau$ gives the equation of motion
$S^{a}_{\;\;;a}=0,$ (7)
with
$S^{a}\equiv{2\over Q}{\cal J}_{,A}A^{a}U^{b}\nabla_{b}\ln{Q}+\Big{[}{2\over Q}({\cal J}_{,A}A^{b})_{;b}-{\cal K}_{,Q}c^{2}\Big{]}U^{a}.$ (8)
Using the equation of motion, we can show $(T^{ab}_{\cal J}+T^{ab}_{\cal K})_{;b}=0$.
We will often consider two models for ${\cal K}$ Blanchet-Skordis-2024
$\hskip-22.76228pt{\rm Model\;I}:\hskip 3.69899pt{\cal K}={2\nu^{2}\over n+1}{\cal K}_{n+1}(Q-1)^{n+1},$ (9)
$\hskip-22.76228pt{\rm Model\;II}:{\cal K}=\sum_{n=1}^{\infty}{2\nu^{2}\over n+1}{\cal K}_{n+1}(Q-1)^{n+1},$ (10)
with ${\cal K}_{2}\equiv 1$.
III Cosmological Background
We consider a flat Friedmann cosmology with the Robertson-Walker metric
$ds^{2}=-a^{2}dx^{0}dx^{0}+a^{2}\delta_{ij}dx^{i}dx^{j},$ (11)
where $x^{0}=\eta$ is the conformal time; $a$ is the cosmic scale factor. Using $\tau=\tau(t)$, we have
$U_{0}=-a,\quad U_{i}=0,\quad U^{0}={1\over a},\quad U^{i}=0;$
$Q=\dot{\tau},\quad A_{a}=0=A,$ (12)
where an overdot indicates a derivative with respect to the cosmic time $t$. Equations (4), with $T_{00}=a^{2}\mu$ and $T_{ij}=a^{2}p\delta_{ij}$, gives $\tau$-fluid quantities
$\mu_{\tau}={c^{4}\over 8\pi G}({\cal J}-{\cal K}+Q{\cal K}_{,Q}),\quad p_{\tau}={c^{4}\over 8\pi G}(-{\cal J}+{\cal K}).$ (13)
As we have $A=0$, only ${\cal J}={\rm constant}$ is allowed which is absorbed to the cosmological constant; thus, we set ${\cal J}=0$.
Einstein’s equation gives
$\hskip-28.45274ptH^{2}={8\pi G\over 3c^{2}}\mu_{m}-{c^{2}\over 3}({\cal K}-Q{\cal K}_{,Q})+{\Lambda c^{2}\over 3},$ (14)
$\hskip-28.45274pt{\ddot{a}\over a}=-{4\pi G\over 3c^{2}}(\mu_{m}+3p_{m})-{c^{2}\over 3}\Big{(}{\cal K}+{1\over 2}Q{\cal K}_{,Q}\Big{)}+{\Lambda c^{2}\over 3},$ (15)
where $\mu_{m}$ and $p_{m}$ are the fluid part of energy density and pressure, respectively; these are sum of radiation, baryon, etc., with respective conservation equations $\dot{\mu}_{m}+3H(\mu_{m}+p_{m})=0$; $H\equiv\dot{a}/a$.
The equation of motion gives $(a^{3}{\cal K}_{,Q})^{\prime}=0$, thus Blanchet-Skordis-2024
${\cal K}_{,Q}={I_{0}\over a^{3}}.$ (16)
We can show
$w_{\tau}\equiv{p_{\tau}\over\mu_{\tau}}={{\cal K}\over-{\cal K}+Q{\cal K}_{,Q}},\quad c_{\tau}^{2}\equiv{\dot{p}_{\tau}\over\dot{\mu}_{\tau}}={{\cal K}_{,Q}\over Q{\cal K}_{,QQ}},$ (17)
and
$\dot{\mu}_{\tau}+3H(\mu_{\tau}+p_{\tau})=0.$ (18)
III.1 Background constraint
For model-I in Eq. (9), Eqs. (16), (13) and (17) give
$\hskip-14.22636ptQ=1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n},$
$\hskip-14.22636pt\varrho_{\tau}={c^{2}I_{0}\over 8\pi Ga^{3}}\Big{[}1+{n\over n+1}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]},$
$\hskip-14.22636ptw_{\tau}=\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{[}n+1+n\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{-1},$
$\hskip-14.22636ptc_{\tau}^{2}={1\over n}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{-1},$ (19)
where $a_{0}=a(t_{0})\equiv 1$ at present epoch; we used $\mu=\varrho c^{2}$. By a suitable parameter choice, with $[I_{0}/(2\nu^{2}{\cal K}_{n+1}a^{3})]^{1/n}\ll 1$, ${\cal K}$-term can simulate a dust fluid with $w_{\tau}=c_{\tau}^{2}=0$ in recent era. In the early era it approaches the stiff fluid with $w_{\tau}=c_{\tau}^{2}=1$ Blanchet-Skordis-2024 .
For the dust-like behavior of $\tau$-field in the background, the authors of Blanchet-Skordis-2024 demand $w_{\tau*}\leq 0.0164$ at $a_{*}\sim 10^{-4.5}$. Using $I_{0}\simeq 8\pi G\varrho_{\tau 0}/c^{2}=3\Omega_{\tau 0}H_{0}^{2}/c^{2}$, we have
${1\over\nu\sqrt{{\cal K}_{n+1}}}\leq\Big{[}{2c^{2}\over 3H_{0}^{2}\Omega_{\tau 0}}\Big{(}{(n+1)w_{\tau*}\over 1-nw_{\tau*}}\Big{)}^{n}a_{*}^{3}\Big{]}^{1/2}\equiv L_{*}.$ (20)
For $n=1$, with $\Omega_{\tau 0}=0.26$, the upper limit gives $L_{*}=0.22kpc$. As $n$ increases the upper limit increases: we have $L_{*}=0.34$, $0.46$, $0.58$, $0.71$, and $1.4kpc$, respectively, for $n=2,3,4,5$, and $10$. This is a constraint on $\nu$-parameter from the requirement of dust-like behavior of $\tau$-field in the background cosmology Blanchet-Skordis-2024 .
IV PN approximation
The 1PN quantities required for Einstein’s equation with a general fluid are presented in Appendix B of Hwang-Noh-axion-2023 . To 1PN order, the metric is
$g_{00}=-\Big{(}1+{2\over c^{2}}\Phi\Big{)},\quad g_{0i}=-a{1\over c^{3}}P_{i},$
$g_{ij}=a^{2}\Big{(}1-{2\over c^{2}}\Psi\Big{)}\delta_{ij},$ (21)
where the index of $P_{i}$ is associated with $\delta_{ij}$ and its inverse. In this form we already imposed spatial gauge conditions Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008 . In PN approximation, we use $x^{0}=ct$. To 0PN order, $\Phi$ is the Newtonian gravitational potential; $\Psi$ and $P_{i}$ are 1PN order. To 1PN order, $\Phi$ in Eq. (59) contains 1PN order contribution as well. In Chandrasekhar’s notation, we have Chandrasekhar-1965 ; Hwang-Noh-Puetzfeld-2008
$\Phi=-U-{1\over c^{2}}\left(2\overline{\Phi}-U^{2}\right),\quad\Psi=-V.$ (22)
For the $\tau$-field, as an ansatz, we expand Flanagan-2023 ; Blanchet-Skordis-2024
$\tau\equiv t+{1\over c^{2}}\sigma+{1\over c^{4}}\Sigma+\dots,$ (23)
where $\sigma$ and $\Sigma$ are functions of spacetime corresponding to 0PN and 1PN order, respectively. As we take $\tau=t$ to the background order, we have $Q=1$ thus ${\cal K}=0$, and the $\tau$-field does not contribute to the background order equations.
From Eq. (2), using the inverse metric in Eq. (B3) of Hwang-Noh-axion-2023 , we have
$\hskip-22.76228ptQ=1+{1\over c^{2}}\Big{(}-\Phi+\dot{\sigma}-{1\over 2a^{2}}\sigma^{,i}\sigma_{,i}\Big{)}$
$\qquad\hskip-22.76228pt+{1\over c^{4}}\Big{\{}{3\over 2}\Phi^{2}+\dot{\Sigma}-\Phi\dot{\sigma}+{1\over a}P^{i}\sigma_{,i}-{1\over a^{2}}\Big{[}{1\over 8a^{2}}(\sigma^{,i}\sigma_{,i})^{2}$
$\qquad\hskip-22.76228pt-{1\over 2}\dot{\sigma}\sigma^{,i}\sigma_{,i}+\sigma^{,i}\Sigma_{,i}+\Big{(}\Psi+{1\over 2}\Phi\Big{)}\sigma^{,i}\sigma_{,i}\Big{]}\Big{\}}.$ (24)
Apparently, the expansion is lengthy if we keep perturbations of $\tau$-field. In the following, we consider $\sigma$ to 00PN order, and for simplicity, set $\sigma=0=\Sigma$ to $1$PN order. The meaning of $\sigma=0=\Sigma$ will be explained later.
IV.1 0PN limit
To 0PN (Newtonian) order, we have
$Q=1-{1\over c^{2}}\Xi,\quad\Xi\equiv\Phi-\dot{\sigma}+{1\over 2a^{2}}\sigma^{,i}\sigma_{,i}.$ (25)
The four-vector and acceleration in Eq. (2) give
$\hskip-28.45274ptU_{0}=-\Big{[}1+{1\over c^{2}}(\Xi+\dot{\sigma})\Big{]},\quad U_{i}=-{1\over c}\sigma_{,i},$
$\hskip-28.45274ptU^{0}=1+{1\over c^{2}}(-2\Phi+\Xi+\dot{\sigma}),\quad U^{i}=-{1\over c}{1\over a^{2}}\sigma^{,i};$
$\hskip-28.45274ptA_{i}=\Xi_{,i},\quad A_{0}={1\over c}{1\over a^{2}}\sigma^{,i}\Xi_{,i},\quad A={1\over c^{4}}{1\over a^{2}}\Xi^{,i}\Xi_{,i}.$ (26)
For $\sigma=0$, we have $U_{a}=n_{a}$.
As a concrete example, we consider the model-II in Eq. (10). We have
${\cal K}={1\over c^{4}}\nu^{2}\Xi^{2},\quad Q{\cal K}_{,Q}=-{2\over c^{2}}\nu^{2}\Xi.$ (27)
Later analysis shows that the nonvanishing ${\cal K}$ is 1PN order. To 0PN order the results are exactly the same as using the model-I in Eq. (9) with $n=1$. That is, only $n=1$ contributes to a mass-like term to 0PN order Blanchet-Skordis-2024 .
The observer’s four-vector $u_{a}$ to 1PN order is presented in Eq. (B7) of Hwang-Noh-axion-2023 ; to 0PN order we have $u_{i}=av_{i}/c$ where the index of $v_{i}$ is associated with $\delta_{ij}$ and its inverse. Using
$q_{i}\equiv{a\over c}Q_{i},\quad\pi_{ij}\equiv a^{2}\Pi_{ij},$ (28)
where indices of $Q_{i}$ and $\Pi_{ij}$ are associated with $\delta_{ij}$ and its inverse, the $\tau$-fluid quantities in Eq. (6) become
$\mu={1\over 8\pi G}\Big{\{}{\cal J}c^{4}-2c^{2}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}\Big{\}},$
$p=-{1\over 8\pi G}{\cal J}c^{4},$
$Q_{i}={2c^{2}\over 8\pi G}\Big{(}v_{i}+{1\over a}\sigma_{,i}\Big{)}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,j})_{,j}+\nu^{2}\Xi\Big{]},$
$\Pi_{ij}=0.$ (29)
To 0PN order, we have $\mu=\varrho c^{2}$. As we set ${\cal J}=0$ to the background order, the $\tau$-fluid quantities do not contribute to the background. As we have ${\cal J}=\delta{\cal J}={\cal J}_{,A}A$ and $A\sim{\cal O}(c^{-4})$, the first term in $\mu$ is higher PN order compared with the second term, and similarly the $p$ is higher PN order compared with $\mu$. Thus, we have
$\varrho=-{1\over 4\pi G}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}.$ (30)
As we have overlapping degrees of freedom in $u_{a}$ (velocity) and $q_{a}$ (flux), we can always take $q_{a}=0$ as a frame choice, thus set $Q_{i}=0$; this is often termed as the energy frame. Under this condition, we have
$v_{i}=-{1\over a}\sigma_{,i}.$ (31)
The $\tau$-fluid is coupled only to the longitudinal part of the velocity. The energy-momentum tensor of $\tau$-field becomes
$T^{0}_{0}=-\mu,\quad T^{0}_{i}=-c\varrho\sigma_{,i},\quad T^{i}_{0}={c\over a^{2}}\varrho\sigma^{,i},\quad T^{i}_{j}=p\delta^{i}_{j}.$ (32)
The energy-momentum conservation, $T^{b}_{a;b}=0$, for the $\tau$-field, gives
${1\over a^{3}}(a^{3}S^{0})^{\displaystyle{\cdot}}-{1\over a^{2}}(S^{0}\sigma^{,i})_{,i}=0,$ (33)
$2\Phi_{,i}\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,j})_{,j}+\nu^{2}\Xi\Big{]}=-{\cal J}_{,i}c^{4},$ (34)
where
$S^{0}=2\Big{[}{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}+\nu^{2}\Xi\Big{]}=-8\pi G\varrho,$ (35)
and $S^{i}=-S^{0}\sigma^{,i}/(ca^{2})$. Equation (33) is the same as equation of motion. In terms of $\tau$-fluid quantities, Eqs. (33) and (34) become
$\dot{\varrho}+3H\varrho-{1\over a^{2}}(\varrho\sigma^{,i})_{,i}=0,$ (36)
$\varrho\Phi_{,i}=-p_{,i}.$ (37)
To 0PN order, using curvature tensors presented in Eq. (B6) of Hwang-Noh-axion-2023 , and using Eq. (32) we can derive Einstein’s equation valid to 0PN order. To the background order, we have Eqs. (14) and (15). By subtracting the background equations, to 0PN order, $R^{i}_{j}$- and $R^{0}_{0}$-components of Einstein’s equation, respectively, give $\Psi=\Phi$ and
${\Delta\over a^{2}}\Phi=4\pi G(\delta\varrho+\delta\varrho_{\tau})$
$\qquad=4\pi G\delta\varrho-{1\over a^{2}}({\cal J}_{,A}\Xi^{,i})_{,i}-\nu^{2}\Xi,$ (38)
where $\delta\varrho$ is the perturbed density of ordinary fluids. This is Poisson’s equation modified by $\tau$-field to 0PN order.
IV.1.1 Gauge transformation to 1PN order
Gauge transformation property in the PN approximation was studied in Sec. 6 of Hwang-Noh-Puetzfeld-2008 . Under a gauge (infinitesimal coordinate) transformation, $\widehat{x}^{a}=x^{a}+\xi^{a}$, we have
$\tau(x^{e})=\widehat{\tau}(\widehat{x}^{e})=\widehat{\tau}(x^{e})+\tau_{,c}\xi^{c}.$ (39)
To 1PN order, as an ansatz, we consider
$\xi^{0}={1\over c}\xi^{(2)0}+{1\over c^{3}}\xi^{(4)0},\quad\xi^{i}={1\over c^{2}}{1\over a}\xi^{(2)i}.$ (40)
Under a spatial gauge condition already imposed in our metric convention in Eq. (21), we have $\xi^{(2)i}=0$, thus $\xi^{i}=0$ Hwang-Noh-Puetzfeld-2008 .
To 0PN and 1PN orders, respectively, we have
$\widehat{\sigma}=\sigma-\xi^{(2)0},\quad\widehat{\Sigma}=\Sigma-\xi^{(4)0}-(\sigma-\xi^{(2)0})\xi^{(2)0}.$ (41)
From the gauge transformation of the metric $g_{0i}$, under our metric ansatz in Eq. (21), we have $\xi^{(2)0}=\xi^{(2)0}(t)$ Hwang-Noh-Puetzfeld-2008 . The spatially constant $\xi^{(2)0}$ can be absorbed by a global redefinition of the time coordinate, and without losing generality we can set $\xi^{(2)0}\equiv 0$; this is necessary to maintain $V=U$, thus $\Psi=\Phi$ to 0PN order Hwang-Noh-Puetzfeld-2008 . Therefore, $\sigma$ cannot be used to fix the temporal gauge condition; in another words, to 0PN order, the PN approximation does not depend on the gauge Hwang-Noh-Puetzfeld-2008 . To 1PN order $\Sigma\equiv 0$ is a legitimate gauge condition which removes the gauge mode completely.
Still, we may set $\sigma=0$ as a physical condition on the $\tau$-fluid Bonetti-2015 . In this case, Eq. (31) implies $v_{i}=0$, thus comoving with the $\tau$-fluid. To 1PN order, we expect $\Sigma=0$ also implies comoving with the $\tau$-fluid, see $T^{0}_{i}$ in Eq. (54). In perturbation theory later, $\sigma$ and $\Sigma$ can be used as the temporal gauge condition which removes the gauge degree of freedom completely; $\sigma\equiv 0\equiv\Sigma$ corresponds to the comoving gauge of the $\tau$-fluid, see Sec. V.1.1. In our PN approximation, however, condition $\sigma=0$ to 0PN order should be considered as a physical condition on the $\tau$-field.
IV.1.2 MOND regime with $\sigma=0$
Equation (38) can be arranged as
${1\over a^{2}}\nabla\cdot\big{(}\nabla\Phi+{\cal J}_{,A}\nabla\Xi\big{)}+\nu^{2}\Xi=4\pi G\delta\varrho.$ (42)
For $\sigma=0$, we have $\Xi=\Phi$, thus
${1\over a^{2}}\nabla\cdot\big{[}(1+{\cal J}_{,A})\nabla\Phi\big{]}+\nu^{2}\Phi=4\pi G\delta\varrho.$ (43)
We have $A_{i}=\Phi_{,i}$, thus the acceleration becomes gravitational acceleration. A non-relativistic Lagrangian formulation of MOND by Bekenstein and Milgrom Bekenstein-Milgrom-1984 , demands
$1+{\cal J}_{,A}=f(|\nabla\Phi|/a_{M})\equiv f(x),$ (44)
with $f(x)=x$ for $x\ll 1$ and $f(x)=1$ for $x\gg 1$. We have
$A={1\over a^{2}c^{4}}\Phi^{,i}\Phi_{,i}={1\over a^{2}c^{4}}|\nabla\Phi|^{2}={1\over a^{2}c^{4}}a_{M}^{2}x^{2}.$ (45)
In order to achieve the MOND and the proper Newtonian gravity in the low and high acceleration regimes, respectively, we demand the following. (i) In the MOND regime, $f(x)=x$ for $x\ll 1$ (MOND regime), we need ${\cal J}_{,A}=x-1=ac^{2}\sqrt{A}/a_{M}-1$, thus Blanchet-Skordis-2024
${\cal J}=-A+{2\over 3}{ac^{2}\over a_{M}}A^{3/2}+{\cal O}(x^{4}).$ (46)
(ii) In the Newtonian regime, we need $f(x)=1$ for $x\gg 1$. This implies ${\cal J}_{,A}$ should be negligibly small compared with unity in that limit Blanchet-Skordis-2024 . For a simple interpolating function $f(x)=x/(1+x)$, we have ${\cal J}_{,A}=-1/(1+x)$ and
${\cal J}=-2{a_{M}\over ac^{2}}\Big{[}\sqrt{A}-{a_{M}\over ac^{2}}\ln{\Big{(}1+{ac^{2}\over a_{M}}\sqrt{A}\Big{)}}\Big{]}.$ (47)
In galactic and smaller scales where MOND phenomena apply the cosmic expansion effect is negligible and we may set $a\equiv 1$. Setting $\tau=t$ for the background is consistent with this. In expanding cosmological background, as in Sec. III, $\tau$-field with $\tau(t)$ contributes to the background evolution through ${\cal K}$. In our PN approximation, as we consider $\tau=t$ for the background, the $\tau$-field does not contribute to the background. For $\tau=\tau(t)$ in the background, it is unclear whether a proper PN expansion is possible. Similarly, the requirement of a proper PN expansion constrains allowed functional form of ${\cal K}$.
The MOND regime naturally appears by choosing ${\cal J}$-function in Eq. (46). Although stationary condition is mentioned in Flanagan-2023 ; Blanchet-Skordis-2024 , we did not impose such a condition. We emphasize that setting $\sigma=0$ is a physical condition imposed on the $\tau$-field.
IV.1.3 MOND constraint for $n=1$
The MOND limit is achieved by the ${\cal J}$-term. The ${\cal K}$-term is introduced only to account for the cosmological demands of the dark matter in the background and perturbations including the large-scale structure and CMB. The ${\cal K}$-term, however, introduces a mass-like correction term in the modified Poisson’s equation, $\nu^{2}\Phi$-term in Eq. (43) with $\nu=m_{\nu}c/\hbar$, which should be negligible. This term appears only for $n=1$ in model-II in Eq. (10). In order to recover phenomenology in galactic scale, we may need Blanchet-Skordis-2024
$\nu^{-1}>1Mpc,$ (48)
corresponding to $m_{\nu}c^{2}<6.4\times 10^{-30}{\rm eV}$. This constraint for a proper MOND phenomenology conflicts with the requirement for dust-like behavior in the background cosmology in Eq. (20), see Blanchet-Skordis-2024 . Thus, $n=1$ part is excluded, and for $n\geq 2$ the model is free from this constraint.
IV.2 1PN approximation with $\sigma=0=\Sigma$
To 1PN order, for simplicity we set $\sigma=0=\Sigma$. From Eq. (24), we have
$Q=1-{1\over c^{2}}\Phi+{3\over 2c^{4}}\Phi^{2}.$ (49)
Equation (2) gives
$U_{0}=-\Big{(}1+{1\over c^{2}}\Phi-{1\over 2c^{4}}\Phi^{2}\Big{)},\quad U_{i}=0,$
$U^{0}=1-{1\over c^{2}}\Phi+{3\over 2c^{4}}\Phi^{2},\quad U^{i}={1\over c^{3}}{1\over a}P^{i};$
$A_{0}=-{1\over c^{3}}{1\over a}P^{i}\Phi_{,i},\quad A_{i}=\Phi_{,i}\Big{(}1-{2\over c^{2}}\Phi\Big{)},$
$A^{0}=0,\quad A^{i}={1\over a^{2}}\Phi^{,i}\Big{[}1+{2\over c^{2}}(\Psi-\Phi)\Big{]},$
$A={1\over c^{4}}{1\over a^{2}}\Phi^{,i}\Phi_{,i}\Big{[}1+{2\over c^{2}}(\Psi-2\Phi)\Big{]}.$ (50)
As we set $\sigma=0=\Sigma$, $U_{a}$ is the same as the normal four-vector. By taking the model-II in Eq. (10), we have
${\cal K}={1\over c^{4}}\nu^{2}\Phi^{2},\quad Q{\cal K}_{,Q}=-{2\over c^{2}}\nu^{2}\Phi\Big{[}1-{1\over c^{2}}\Big{(}{5\over 2}+{\cal K}_{3}\Big{)}\Phi\Big{]}.$ (51)
Thus, only $n=1$ and $2$ contribute to 1PN order.
Using Eqs. (4), (6) and (28), the $\tau$-fluid quantities are
$\hskip-14.22636pt\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}+v^{2}+2\Psi-2\Phi)+c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}+v^{2}-(2+{\cal K}_{3})\Phi\big{]}\Big{\}},$
$\hskip-14.22636ptp={1\over 8\pi G}{1\over 3}\Big{[}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}v^{2}-3c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}-\nu^{2}\Phi(2v^{2}-3\Phi)\Big{]},$
$\hskip-14.22636ptQ_{i}={1\over 8\pi G}\Big{\{}{2\over a^{2}}v_{i}({\cal J}_{,A}\Phi^{,j})_{,j}\Big{(}c^{2}+{3\over 2}v^{2}+2\Psi-2\Phi\Big{)}-{2\over a^{2}}{\cal J}_{,A}\big{[}\Phi_{,i}\Phi^{,j}v_{j}+v_{i}(\Phi+\Psi)^{,j}\Phi_{,j}\big{]}$
$\qquad\hskip-14.22636pt+\nu^{2}\Phi v_{i}\big{[}2c^{2}+3v^{2}-(5+2{\cal K}_{3})\Phi\big{]}\Big{\}},$
$\hskip-14.22636pt\Pi_{ij}={1\over 8\pi G}\Big{[}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,k})_{,k}\Big{(}v_{i}v_{j}-{1\over 3}v^{2}\delta_{ij}\Big{)}+{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\Phi^{,k}\Phi_{,k}\delta_{ij}\Big{)}-2\nu^{2}\Phi\Big{(}v_{i}v_{j}-{1\over 3}v^{2}\delta_{ij}\Big{)}\Big{]}.$ (52)
To 1PN order, we have $\mu=\varrho(c^{2}+\Pi)$ with $\varrho\Pi$ the internal energy density Chandrasekhar-1965 . The energy-frame condition $q_{i}\equiv 0$ for the $\tau$-fluid implies $v_{i}=0$, thus the velocity four-vector becomes normal, $u_{a}=n_{a}$. The fluid quantities become
$\mu={1\over 8\pi G}\Big{\{}-{2\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}(c^{2}+2\Psi-2\Phi)+c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}(\Phi+\Psi)^{,i}\Phi_{,i}-2\nu^{2}\Phi\big{[}c^{2}-(2+{\cal K}_{3})\Phi\big{]}\Big{\}},$
$p={1\over 8\pi G}{1\over 3}\Big{(}-3c^{4}{\cal J}+{2\over a^{2}}{\cal J}_{,A}\Phi^{,i}\Phi_{,i}+3\nu^{2}\Phi^{2}\Big{)},\quad\Pi_{ij}={1\over 8\pi G}{2\over a^{2}}{\cal J}_{,A}\Big{(}\Phi_{,i}\Phi_{,j}-{1\over 3}\Phi^{,k}\Phi_{,k}\delta_{ij}\Big{)}.$ (53)
For the $\tau$-fluid, using Eq. (53), and Eq. (B9) in Hwang-Noh-axion-2023 , we can show
$T^{0}_{0}=-\mu,\quad T^{0}_{i}=0,\quad T^{i}_{0}=-{1\over c^{3}a}\mu P^{i},\quad T^{i}_{j}=p\delta^{i}_{j}+\Big{(}1+{2\over c^{2}}\Psi\Big{)}\Pi^{i}_{j}.$ (54)
We note that setting $q_{i}=0$ is a frame choice always allowed for any energy-momentum component. Under this choice, the flux or momentum of the component is encoded in the four-vector $u_{a}$. As the four-vector becomes normal with $n_{i}\equiv 0$, we have the physical momentum of the component vanish; this is reflected in $T^{0}_{i}=(\mu+p)u^{0}u_{i}=0$ to 1PN order. This implies that the $\tau$-fluid quantities are measured by an observer comoving with the $\tau$-fluid. This happened as we imposed $\sigma=0=\Sigma$ as a physical condition.
Using Eq. (54), $T^{b}_{a;b}=0$ for $\tau$-fluid gives
$\dot{\mu}+3H(\mu+p)-3\varrho\dot{\Psi}+{1\over a}(\varrho P^{i})_{,i}=0,$ (55)
$\varrho\Phi_{,i}+p_{,i}+\Pi^{j}_{i,j}+{1\over c^{2}}\big{\{}\Phi_{,i}[\varrho(\Pi-2\Phi)+p]+(\Phi-\Psi)_{,j}\Pi^{j}_{i}+2\Psi\Pi^{j}_{i,j}\big{\}}=0.$ (56)
These coincide with the energy and momentum conservation equations of a fluid in (B23) and (B24) of Hwang-Noh-axion-2023 by setting $v_{i}=0=Q_{i}$. For a fluid, $Q_{i}=0$ can be regarded as the frame choice, and $v_{i}=0$ implies that the fluid velocity vanishes which can be achieved in relativistic perturbation theory by a temporal gauge choice; the comoving gauge condition for a fluid imposes only the longitudinal part of $v_{i}$ to vanish. In our $\tau$-fluid case, $\sigma=0=\Sigma$ implies $v_{i}=0$, thus the transverse part also vanishes; in the perturbation theory in next section we will show that the $\tau$-field does not support the (transverse) vector-type perturbation.
Equation of motion in Eqs. (7) and (8) gives
$\dot{S}^{0}+\Big{[}3H+{1\over c^{2}}(\dot{\Phi}-3\dot{\Psi})\Big{]}S^{0}+cS^{i}_{\;\;,i}=0,$ (57)
$S^{0}={1\over c^{2}}(-8\pi G\mu+c^{4}{\cal J}-\mu^{2}\Phi^{2}),\quad S^{i}=-{1\over c^{3}}\Big{(}{2\over a^{2}}{\cal J}_{,A}\dot{\Phi}\Phi^{,i}+8\pi G\varrho{1\over a}P^{i}\Big{)},$ (58)
and it gives Eq. (55).
Using the curvature tensors presented in Eq. (B6) of Hwang-Noh-axion-2023 , we can derive Einstein’s equation valid to 1PN order. To the background order, we have Eqs. (14) and (15). By subtracting the background equations, to 1PN order, $R^{0}_{0}$-, $R^{i}_{j}$- and $R^{0}_{i}$-components of Einstein’s equation, respectively, give
$\hskip-19.91684pt{\Delta\over a^{2}}\Phi+{1\over c^{2}}\Big{\{}3\Big{[}\ddot{\Psi}+H(\dot{\Phi}+2\dot{\Psi})+2{\ddot{a}\over a}\Phi\Big{]}-{1\over a^{2}}\big{[}2(\Phi-\Psi)\Delta\Phi+(\Phi+\Psi)^{,i}\Phi_{,i}\big{]}-{1\over a^{2}}(aP^{i}_{\;\;,i})^{\displaystyle{\cdot}}\Big{\}}+\nu^{2}\Phi\Big{[}1-{1\over c^{2}}\Big{(}{7\over 2}+{\cal K}_{3}\Big{)}\Phi\Big{]}$
$\qquad\hskip-19.91684pt=4\pi G\Big{\{}\delta\Big{[}\varrho\Big{(}1+{1\over c^{2}}\Pi\Big{)}\Big{]}+{3\over c^{2}}\delta p+{2\over c^{2}}\varrho v^{2}\Big{\}}-{1\over a^{2}}({\cal J}_{,A}\Phi^{,i})_{,i}\Big{[}1+{2\over c^{2}}(\Psi-\Phi)\Big{]}-{\cal J}c^{2}+{1\over a^{2}c^{2}}{\cal J}_{,A}(2\Phi+\Psi)_{,i}\Phi^{,i},$ (59)
$\hskip-19.91684pt{\Delta\over a^{2}}\Psi\delta^{i}_{j}+{1\over a^{2}}(\Psi-\Phi)^{,i}_{\;\;\;j}=\Big{[}4\pi G\delta\varrho-{1\over a^{2}}({\cal J}_{,A}\Phi^{,k})_{,k}-\nu^{2}\Phi\Big{]}\delta^{i}_{j},$ (60)
$\hskip-19.91684pt-2(\dot{\Psi}+H\Phi)_{,i}+{1\over 2a}(P^{j}_{\;\;,ji}-\Delta P_{i})=8\pi G\varrho av_{i}.$ (61)
Equation (60) is valid to 0PN order whereas the other two equations are valid to 1PN order. To 0PN order, Eq. (59) gives Eq. (38). To 0PN order, Eq. (60) is the same as Eq. (59), with $\Psi=\Phi$.
The fluid quantities in Eqs. (59)-(61) are the sum of other fluids besides the $\tau$-fluid; we may include the $\tau$-fluid to the collective fluid quantities as $\mu=\mu_{m}+\mu_{\tau}$, $p=p_{m}+p_{\tau}$, and $\varrho v_{i}=\varrho^{m}v^{m}_{i}+\varrho^{\tau}v^{\tau}_{i}$. For the $\tau$-fluid we have conservation equations in (55)-(56) or the equation of motion in Eq. (57). The fluid conservation equations to 1PN order are in Eqs. (B23) and (B24) of Hwang-Noh-axion-2023 . These are the complete set of equations valid to 1PN order without imposing the temporal gauge condition; for the choice of gauge conditions, see Sec. 6 in Hwang-Noh-Puetzfeld-2008 .
V Perturbation theory
We consider a metric
$g_{00}=-a^{2}(1+2\alpha),\quad g_{0i}=-a\chi_{i},$
$g_{ij}=a^{2}(1+2\varphi)\delta_{ij},$ (62)
where $\alpha$, $\varphi$ and $\chi_{i}$ are functions of spacetime with arbitrary amplitudes and the index of $\chi_{i}$ is associated with $\delta_{ij}$ and its inverse; here, we have $x^{0}=\eta$. In this form, we ignored the tracefree-transverse tensor type perturbation, and imposed a spatial gauge condition which (together with appropriate choice of temporal gauge condition) removes the gauge degrees of freedom completely. In Hwang-Noh-2013 we derived the fully nonlinear and exact perturbation equations, and the case is extended to more general metric including the tensor type peturbation and without imposing spatial gauge condition in Gong-2017 . The fully nonlinear and exact formulation useful for our analysis is summarized in Appendix C of Hwang-Noh-EM-2023 .
We set
$\tau\equiv\bar{\tau}(t)+{1\over c^{2}}\sigma({\bf x},t),$ (63)
where $\sigma$ is a function of spacetime with arbitrary amplitude; in PN expansion $\Sigma$ and other higher order expansions in Eq. (23) are absorbed to $\sigma$; if needed we use an overbar to indicate the background order quantity. The fully nonlinear and exact perturbation formulation includes the PN expansion; comparing Eqs. (62) and (21) we have
$\alpha={\Phi\over c^{2}},\quad\varphi=-{\Psi\over c^{2}},\quad\chi_{i}=a{P_{i}\over c^{3}}.$ (64)
Equation (2), using the exact inverse metric in Eq. (C3) of Hwang-Noh-EM-2023 , we have
$\hskip-25.6073ptQ={1\over{\cal N}}\Big{[}\Big{(}\dot{\bar{\tau}}+{1\over c^{2}}\dot{\sigma}\Big{)}^{2}+{2\chi^{i}\over a^{2}(1+2\varphi)}\Big{(}\dot{\bar{\tau}}+{1\over c^{2}}\dot{\sigma}\Big{)}{1\over c}\sigma_{,i}$
$\qquad\hskip-25.6073pt-{1\over a^{2}(1+2\varphi)}\Big{(}{\cal N}^{2}\delta^{ij}-{\chi^{i}\chi^{j}\over a^{2}(1+2\varphi)}\Big{)}{1\over c^{2}}\sigma_{,i}\sigma_{,j}\Big{]}^{1/2},$ (65)
with ${\cal N}$ given in Eq. (95). Apparently, the expansion is lengthy if we keep perturbations of $\sigma$. In the following, we consider $\sigma$ in the linear order perturbation theory, and for simplicity, set $\sigma=0$ in the fully nonlinear and exact formulation; $\sigma=0$ corresponds to a temporal gauge condition valid to fully nonlinear order and it is the same as the comoving gauge of the $\tau$-fluid, see Eqs. (72) and (99).
V.1 Linear perturbation
To the linear order in perturbations, Eq. (65) gives
$Q=(1-\alpha)\bar{Q}+{1\over c^{2}}\dot{\sigma},\quad\bar{Q}=\dot{\bar{\tau}}.$ (66)
The four-vector and acceleration become
$U_{0}=-a(1+\alpha),\quad U_{i}=-{1\over cQ}\sigma_{,i},$
$U^{0}={1\over a}(1-\alpha),\quad U^{i}={1\over a^{2}}\Big{(}\chi^{i}-{1\over cQ}\sigma^{,i}\Big{)};$
$A_{0}=0,\quad A_{i}=c^{2}\Upsilon_{,i},\quad A^{0}=0,\quad A^{i}={c^{2}\over a^{2}}\Upsilon^{,i},$
$\Upsilon\equiv\alpha-{1\over c^{2}Q}\Big{(}\dot{\sigma}-{\dot{Q}\over Q}\sigma\Big{)}\equiv\alpha_{\sigma},\quad A=0,$ (67)
where $\alpha_{\sigma}$ is a unique gauge-invariant notation which becomes $\alpha$ in the $\sigma=0$ gauge condition, see later. The $\tau$-field energy-momentum tensor becomes
$\hskip-14.22636ptT_{00}={c^{4}\over 8\pi G}a^{2}\Big{[}(1+2\alpha)({\cal J}-{\cal K}+Q{\cal K}_{,Q})-2{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{]},$
$\hskip-14.22636ptT_{0i}={c^{4}\over 8\pi G}a\Big{[}({\cal J}-{\cal K})\chi_{i}+{\cal K}_{,Q}{1\over c}\sigma_{,i}\Big{]},$
$\hskip-14.22636ptT_{ij}={c^{4}\over 8\pi G}a^{2}(1+2\varphi)(-{\cal J}+{\cal K})\delta_{ij}.$ (68)
The ADM fluid quantities are
$E\equiv n_{a}n_{b}T^{ab},\quad J_{i}\equiv-n_{b}T^{b}_{i},\quad S_{ij}\equiv T_{ij}.$ (69)
where indices of $J_{i}$ and $S_{ij}$ are associated with the intrinsic ADM metric, $\overline{h}_{ij}(\equiv g_{ij})$, and its inverse. Using Eqs. (117) and (C1) of Hwang-Noh-EM-2023 we have
$E\equiv\mu,\quad J_{i}\equiv acm_{i},\quad S_{ij}\equiv a^{2}(1+2\varphi)m_{ij},$ (70)
where indices of $m_{i}$ and $m_{ij}$ are associated with $\delta_{ij}$ and its inverse. Using
$m_{i}={1\over c^{2}}(\mu+p)v_{i},\quad v_{i}\equiv-v_{,i}+v^{(v)}_{i},\quad v^{(v)i}_{\;\;\;\;\;\;,i}\equiv 0,$
$p={1\over 3}m^{k}_{k},\quad\overline{m}_{ij}\equiv m_{ij}-{1\over 3}m^{k}_{k}\delta_{ij},$ (71)
we have the $\tau$-fluid quantities
$\mu={c^{4}\over 8\pi G}\Big{(}{\cal J}-{\cal K}+Q{\cal K}_{,Q}-2{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{)},$
$(\mu+p)v={c^{4}\over 8\pi G}{1\over a}{\cal K}_{,Q}\sigma,\quad v^{(v)}_{i}=0,$
$p={c^{4}\over 8\pi G}(-{\cal J}+{\cal K}),\quad\overline{m}_{ij}=0.$ (72)
To the background order, we recover Eq. (13), and only scalar-type perturbations are excited without anisotropic stress. We have
${\delta p\over\delta\mu}=c_{\tau}^{2}\Big{(}1-{2a^{3}c_{\tau}^{2}\over I_{0}\delta Q}{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{)}^{-1}.$ (73)
The entropic perturbation of $\tau$-fluid is
$e\equiv\delta p-c_{\tau}^{2}\delta\mu={c^{4}\over 8\pi G}2c_{\tau}^{2}{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon.$ (74)
As we have $A=0$ and ignore the constant part of ${\cal J}$, we have ${\cal J}=0$ whereas ${\cal J}_{,A}\neq 0$. We note that for vanishing ${\cal K}$, we have $p_{\tau}=0=c_{\tau}^{2}$, thus $e=0$.
Einstein’s equations in Eqs. (C8)-(C12) of Hwang-Noh-EM-2023 give
$\hskip-22.76228pt\kappa\equiv 3H\alpha-3\dot{\varphi}-c{\Delta\over a^{2}}\chi,$ (75)
$\hskip-22.76228ptH\kappa+c^{2}{\Delta\over a^{2}}\varphi=-{4\pi G\over c^{2}}\delta\mu,$ (76)
$\hskip-22.76228pt\kappa+c{\Delta\over a^{2}}\chi={12\pi G\over c^{4}}a(\mu+p)v,$ (77)
$\hskip-22.76228pt\dot{\kappa}+2H\kappa+\Big{(}c^{2}{\Delta\over a^{2}}+3\dot{H}\Big{)}\alpha=-{4\pi G\over 3c^{2}}(\delta\mu+\delta p),$ (78)
$\hskip-22.76228ptc(\dot{\chi}+H\chi)-c^{2}(\alpha+\varphi)=0,$ (79)
where the fluid quantities are the collective ones, for example, $\delta\mu=\delta\mu_{m}+\delta\mu_{\tau}$, $\delta p=\delta p_{m}+\delta p_{\tau}$, $(\mu+p)v=(\mu_{m}+p_{m})v_{m}+(\mu_{\tau}+p_{\tau})v_{\tau}$; we decomposed $\chi_{i}\equiv\chi_{,i}+\chi_{i}^{(v)}$ with $\chi^{(v)i}_{\;\;\;\;\;\;,i}\equiv 0$. We ignored the anisotropic stress of ordinary fluid in Eq. (79). The vector- and tensor-type perturbation equations are not affected by the $\tau$-field.
The equation of motion in Eq. (7) gives
${\Delta\over a}(a{\cal J}_{,A}\Upsilon)^{\displaystyle{\cdot}}={\bar{Q}\over 2a}\big{\{}a^{3}\big{[}(1+3\varphi)\bar{\cal K}_{,Q}+\delta({\cal K}_{,Q})\big{]}\big{\}}^{\displaystyle{\cdot}}$
$\qquad+{c\over 2}Q{\cal K}_{,Q}\Delta\Big{(}\chi-{1\over cQ}\sigma\Big{)}.$ (80)
To background order, we recover Eq. (16).
The energy and momentum conservation equations in Eqs. (C14) of Hwang-Noh-EM-2023 give
$\hskip-28.45274pt\delta\dot{\mu}+3H(\delta\mu+\delta p)=(\mu+p)\Big{(}\kappa-3H\alpha+{\Delta\over a}v\Big{)},$ (81)
$\hskip-28.45274pt{1\over a^{4}}[a^{4}(\mu+p)v]^{\displaystyle{\cdot}}={c^{2}\over a}[(\mu+p)\alpha+\delta p].$ (82)
These two equations are valid for ordinary fluid with the anisotropic stress ignored in the second equation, and are also valid for the $\tau$-fluid with the $\tau$-fluid quantities given in Eq. (72). Using the $\tau$-fluid quantities, Eq. (81) is naturally valid by the equation of motion in Eq. (80) together with $\dot{\cal J}=0$. Using Eqs. (16), (66) and (72), Eq. (82) is identically valid.
The complete set of linear perturbation equations without imposing temporal gauge condition is Eqs. (75)-(79) from Einstein’s equation, Eqs. (81) and (82) for other fluids, and the same equation for the $\tau$-fluid with the equation of state in Eqs. (72) and (74). For the $\tau$-fluid, instead of conservation equations in (81) and (82), we can directly use the field variable and the equation of motion in Eq. (80).
V.1.1 Gauge transformation
Under the gauge transformation, $\widehat{x}^{a}=x^{a}+\xi^{a}$, using Eq. (39), Eqs. (63) and (66) give
${\widehat{\sigma}\over c^{2}}={\sigma\over c^{2}}-\tau^{\prime}\xi^{0},\quad\widehat{Q}=Q-c\Big{(}{\tau^{\prime}\over a}\Big{)}^{\prime}\xi^{0},$
$\widehat{\cal K}={\cal K}-{\cal K}_{,Q}c\Big{(}{\tau^{\prime}\over a}\Big{)}^{\prime}\xi^{0},$ (83)
and $\Upsilon$, denoted as $\alpha_{\sigma}$, is gauge-invariant, thus ${\cal J}$ is also gauge-invariant. For $\tau$-fluid quantities, we can show
$\widehat{\mu}=\mu-\mu^{\prime}\xi^{0},\quad\widehat{p}=p-p^{\prime}\xi^{0},\quad\widehat{v}=v-c\xi^{0},$ (84)
which are also valid for ordinary fluids. For gauge transformation properties of the metric, see Eq. (250) of Noh-Hwang-2004 .
From Eq. (83) we notice that $\sigma$ can be used to fix the gauge condition, and $\sigma=0$ in all coordinates gives $\xi^{0}=0$, thus fixes the gauge (coordinate) degree of freedom completely; this implies that under such a gauge condition the remaining perturbation variables are gauge-invariant. Equation (72) shows that $\sigma=0$ implies $v=0$, thus corresponds to the comoving gauge of the $\tau$-fluid. This analysis can be extended to fully nonlinear order, see Sec. VI of Noh-Hwang-2004 .
As a temporal gauge condition we can use $\sigma=0$ in all coordinate systems, and we may call it uniform-$\tau$ gauge. This implies $v=0$ for the $\tau$-fluid, and we may call it the comoving gauge of the $\tau$-fluid. Similarly we can impose $v=0$ for an ordinary fluid. As another choice we can set $\delta\mu=0$ in all coordinates which is the uniform-density gauge for the $\tau$-fluid or ordinary fluid. Other temporal gauge conditions which remove the gauge degree of freedom completely are the following: $\varphi\equiv 0$ (uniform-curvature gauge), $\kappa\equiv 0$ (uniform expansion gauge), $\chi\equiv 0$ (zero-shear gauge) Bardeen-1988 . One exception where the temporal gauge degree of freedom is not completely fixed even after imposing the gauge condition is $\alpha\equiv 0$ (synchronous gauge). These gauge conditions are applicable to nonlinear and exact order with corresponding gauge invariance for other than synchronous gauge Noh-Hwang-2004 ; Hwang-Noh-2013 .
V.1.2 Density perturbation in $\sigma=0$ gauge
We consider density perturbation equation of the $\tau$-field in the $\sigma=0$ gauge condition. The $\sigma=0$ gauge condition implies the comoving gauge ($v=0$) of the $\tau$-fluid, which is true even to fully nonlinear order in perturbation, see Eq. (99). Under this gauge Eqs. (82) and (81), respectively, give
$\alpha_{\sigma}=-{\delta p_{\sigma}\over\mu+p},\quad\kappa_{\sigma}={(a^{3}\delta\mu_{\sigma})^{\displaystyle{\cdot}}\over a^{3}(\mu+p)},$ (85)
where we decomposed $\mu=\bar{\mu}+\delta\mu$ and $p=\bar{p}+\delta p$. A subindex $\sigma$ indicates the $\sigma=0$ gauge imposed, or equivalently a unique gauge-invariant combination; from Eqs. (83) and (84), we have
$\delta\mu_{\sigma}\equiv\delta\mu-{\mu^{\prime}\over\tau^{\prime}}{\sigma\over c^{2}},$ (86)
which is a unique gauge-invariant combination of $\delta\mu$ and $\sigma$ and becomes $\delta\mu$ in the $\sigma=0$ gauge; as long as a gauge condition completely fixes the gauge degree of freedom, all variables in that gauge have unique corresponding gauge-invariant combinations. Ignoring other fluids, Eq. (78) gives
${1+w\over a^{2}H}\Big{[}{H^{2}\over a(\mu+p)}\Big{(}{a^{3}\mu\over H}\delta_{\sigma}\Big{)}^{\displaystyle{\cdot}}\Big{]}^{\displaystyle{\cdot}}=c^{2}{\Delta\over a^{2}}{\delta p_{\sigma}\over\mu},$ (87)
where $\delta\equiv\delta\mu/\mu$; it is valid in the presence of the cosmological constant. The same equation is valid for a general fluid in the absence of anisotropic stress HN-GRG-1999 .
The ${\cal K}(Q)$-term is introduced in the Lagrangian in order to reproduce cosmology which demands presence of dark matter in the context of CMB and the large-scale structure. Here, we are curious whether the pressure terms in the above equation can be ignored in the perturbation level so that the $\tau$-field in the form of ${\cal K}$-term can supplement the large-scale dark matter demanded in current cosmology.
If we can ignore background pressure compared with background energy density, as in the model-I considered in Eq. (20), Eq. (87) gives
$\hskip-17.07182pt{1\over a^{2}H}\Big{[}a^{2}H^{2}\Big{(}{\delta_{\sigma}\over H}\Big{)}^{\displaystyle{\cdot}}\Big{]}^{\displaystyle{\cdot}}=\ddot{\delta}_{\sigma}+2H\dot{\delta}_{\sigma}-4\pi G\varrho\delta_{\sigma}=c^{2}{\Delta\over a^{2}}{\delta p_{\sigma}\over\mu}.$
Comparison of the pressure term with gravity term gives the Jeans scale. Setting $\Delta=-k^{2}$, we have the Jeans wavenumber
${k_{J}\over a}=\sqrt{4\pi G\delta\varrho_{\sigma}\over c^{2}\delta p_{\sigma}/\mu}=\Big{[}2c_{\tau}^{2}\Big{(}{w_{\tau}\over{\cal K}}+{a^{3}\over I_{0}}{1\over Q}{\cal J}_{,A}\Big{)}\Big{]}^{-1/2},$ (89)
where the right-hand side is evaluated in the background; we used Eqs. (16), (17), (72), (66) and (67). The density perturbation of $\tau$-field is stable (unstable) for $k>k_{J}$ ($k<k_{J}$), and it behaves as a cold dark matter for negligible presure perturbation with $k\ll k_{J}$. The above Jeans criterion is valid for general ${\cal J}$ and ${\cal K}$, assuming the background $\tau$-field behaves as a dust fluid.
V.1.3 Perturbation constraint
For the model-I in Eq. (9), we have the background equation of state in Eq. (19). For perturbed part, using Eq. (72), we have
$\delta\mu_{\sigma}=-{c^{4}\over 8\pi G}\Big{\{}n{I_{0}\over a^{3}}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{-1/n}$
×$\qquad\times\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{2}\alpha+2{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{\}},$
$\delta p_{\sigma}=-{c^{4}\over 8\pi G}{I_{0}\over a^{3}}\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}\alpha.$ (90)
For a dust-like equation of state for the background, we have $[I_{0}/(\nu^{2}{\cal K}_{n+1}a^{3})]^{1/n}(\ll 1)$ and Eq. (88) is valid.
The Jeans scale becomes
${k_{J}\over a}\simeq\Big{[}{2\over n}{a^{3}\over I_{0}}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}(1+{\cal J}_{,A})\Big{]}^{-1/2}.$ (91)
For $n=1$, we have
${k_{J}\over a}\simeq{\nu\over\sqrt{1+{\cal J}_{,A}}},$ (92)
thus, the Jeans scale corresponds to the length scale of the mass-like term. In Eq. (48) a reasonable phenomenology in galactic scale demanded $1/\nu>1{\rm Mpc}$ and this applies only for $n=1$. In Eq. (20) a dust-like behavior in the background demands $1/\nu\leq 0.22kpc$. Thus, $n=1$ case is excluded Blanchet-Skordis-2024 .
For $n\geq 2$, we have no constraint in Eq. (48). For example, for $n=2$, we have
${k_{J}\over a}\simeq\sqrt{\sqrt{2I_{0}\over a^{3}}{\nu\sqrt{{\cal K}_{3}}\over 1+{\cal J}_{,A}}}.$ (93)
We can estimate
$L_{J}\leq\sqrt{{L_{H}L_{*}\over\sqrt{6\Omega_{\tau 0}a}}(1+{\cal J}_{,A})}\sim\sqrt{{1+{\cal J}_{,A}\over\sqrt{a}}}1.1Mpc,$ (94)
where we used $k_{J}\sim L_{J}^{-1}$, $L_{H}\equiv c/H_{0}=4.3Gpc$ and Eq. (20). In non-MOND regime ${\cal J}_{,A}$ should be negligible, see Eq. (46) and below.
V.2 Fully nonlinear and exact perturbations in $\sigma=0$ gauge
Setting $\sigma\equiv 0$ as the temporal gauge condition, Eq. (65) significantly simplifies to become
$Q={1\over{\cal N}}\dot{\tau}^{2},\quad{\cal N}\equiv\sqrt{1+2\alpha+{\chi^{k}\chi_{k}\over a^{2}(1+2\varphi)}},$ (95)
where $\tau=\bar{\tau}(t)$. The four-vector becomes
$U_{0}=-a{\cal N},\quad U_{i}=0,$
$U^{0}={1\over a{\cal N}},\quad U^{i}={\chi^{i}\over a^{2}{\cal N}(1+2\varphi)}.$ (96)
This is the same as the normal four-vector in Eq. (93) of Hwang-Noh-EM-2023 . The acceleration vector becomes
$A_{0}={c^{2}\chi^{i}\over a(1+2\varphi)}{Q_{,i}\over Q},\quad A_{i}=-c^{2}{Q_{,i}\over Q},\quad A^{0}=0,\quad A^{i}=-{c^{2}\over a^{2}(1+2\varphi)}{Q^{,i}\over Q},\quad A={Q^{,i}Q_{,i}\over a^{2}(1+2\varphi)Q^{2}}.$ (97)
The energy-momentum tensor of the $\tau$-fluid becomes
$T_{00}={c^{4}\over 8\pi G}\Big{[}a^{2}(1+2\alpha)({\cal J}-{\cal K})+2{\cal J}_{,A}\Big{(}{\chi^{i}Q_{,i}\over a(1+2\varphi)Q}\Big{)}^{2}+{2{\cal N}\over(1+2\varphi)^{3/2}}\Big{(}{\cal N}\sqrt{1+2\varphi}{\cal J}_{,A}{Q^{,i}\over Q}\Big{)}_{,i}+a^{2}{\cal N}^{2}Q{\cal K}_{,Q}\Big{]},$
$T_{0i}={c^{4}\over 8\pi G}\Big{[}({\cal J}-{\cal K})a\chi_{i}-2{\cal J}_{,A}{\chi^{j}Q_{,j}Q_{,i}\over a(1+2\varphi)Q^{2}}\Big{]},\quad T_{ij}={c^{4}\over 8\pi G}\Big{[}-a^{2}(1+2\varphi)({\cal J}-{\cal K})\delta_{ij}+2{\cal J}_{,A}{Q_{,i}Q_{,j}\over Q^{2}}\Big{]}.$ (98)
Using Eqs. (70) and (71), the $\tau$-fluid quantities become
$E={c^{4}\over 8\pi G}\Big{[}{\cal J}-{\cal K}+Q{\cal K}_{,Q}+{2\over{\cal N}(1+2\varphi)^{3/2}}\Big{(}{\cal N}\sqrt{1+2\varphi}{\cal J}_{,A}{Q^{,i}\over Q}\Big{)}_{,i}\Big{]},\quad m_{i}=0,$
$m_{ij}={c^{4}\over 8\pi G}\Big{[}-({\cal J}-{\cal K})\delta_{ij}+{2{\cal J}_{,A}\over a^{2}(1+2\varphi)}{Q_{,i}Q_{,j}\over Q^{2}}\Big{]},\quad\overline{m}_{ij}={c^{4}\over 8\pi G}{2{\cal J}_{,A}\over a^{2}(1+2\varphi)Q^{2}}\Big{(}Q_{,i}Q_{,j}-{1\over 3}Q^{,k}Q_{,k}\delta_{ij}\Big{)}.$ (99)
$m_{i}=0$ shows that for $\sigma=0$ gauge condition the velocity of the $\tau$-fluid vanishes, thus comoving to fully nonlinear order.
Einstein’s equation and the conservation equations in the fully nonlinear and exact formulation with the above fluid quantities are presented in Eqs. (C8)-(C12), (C14) and (C15) of Hwang-Noh-EM-2023 . Einstein’s equations give
$-{3\over 2}\Big{(}{\dot{a}^{2}\over a^{2}}-{8\pi G\over 3c^{2}}E-{\Lambda c^{2}\over 3}\Big{)}+{\dot{a}\over a}\kappa+{c^{2}\Delta\varphi\over a^{2}(1+2\varphi)^{2}}={1\over 6}\kappa^{2}+{3\over 2}{c^{2}\varphi^{,i}\varphi_{,i}\over a^{2}(1+2\varphi)^{3}}-{c^{2}\over 4}\overline{K}^{i}_{j}\overline{K}^{j}_{i},$ (100)
${2\over 3}{1\over c}\kappa_{,i}+{1\over a^{2}{\cal N}(1+2\varphi)}\Big{(}{1\over 2}\Delta\chi_{i}+{1\over 6}\chi^{k}_{\;\;,ki}\Big{)}={1\over a^{2}{\cal N}(1+2\varphi)}\Big{\{}\Big{(}{{\cal N}_{,j}\over{\cal N}}-{\varphi_{,j}\over 1+2\varphi}\Big{)}\Big{[}{1\over 2}\Big{(}\chi^{j}_{\;\;,i}+\chi_{i}^{\;,j}\Big{)}-{1\over 3}\delta^{j}_{i}\chi^{k}_{\;\;,k}\Big{]}$
$\qquad-{\varphi^{,j}\over(1+2\varphi)^{2}}\Big{(}\chi_{i}\varphi_{,j}+{1\over 3}\chi_{j}\varphi_{,i}\Big{)}+{{\cal N}\over 1+2\varphi}\nabla_{j}\Big{[}{1\over{\cal N}}\Big{(}\chi^{j}\varphi_{,i}+\chi_{i}\varphi^{,j}-{2\over 3}\delta^{j}_{i}\chi^{k}\varphi_{,k}\Big{)}\Big{]}\Big{\}}-{8\pi G\over c^{3}}am_{i},$ (101)
$-3\Big{[}{1\over{\cal N}}\Big{(}{\dot{a}\over a}\Big{)}^{\displaystyle\cdot}+{\dot{a}^{2}\over a^{2}}+{4\pi G\over 3c^{2}}\Big{(}E+S\Big{)}-{\Lambda c^{2}\over 3}\Big{]}+{1\over{\cal N}}\dot{\kappa}+2{\dot{a}\over a}\kappa+{c^{2}\Delta{\cal N}\over a^{2}{\cal N}(1+2\varphi)}$
$\qquad={1\over 3}\kappa^{2}-{c\over a^{2}{\cal N}(1+2\varphi)}\Big{(}\chi^{i}\kappa_{,i}+c{\varphi^{,i}{\cal N}_{,i}\over 1+2\varphi}\Big{)}+c^{2}\overline{K}^{i}_{j}\overline{K}^{j}_{i},$ (102)
$\Big{(}{1\over{\cal N}}{\partial\over\partial t}+3{\dot{a}\over a}-\kappa+{c\chi^{k}\over a^{2}{\cal N}(1+2\varphi)}\nabla_{k}\Big{)}\Big{\{}{c\over a^{2}{\cal N}(1+2\varphi)}$
×$\qquad\times\Big{[}{1\over 2}\Big{(}\chi^{i}_{\;\;,j}+\chi_{j}^{\;\;,i}\Big{)}-{1\over 3}\delta^{i}_{j}\chi^{k}_{\;\;,k}-{1\over 1+2\varphi}\Big{(}\chi^{i}\varphi_{,j}+\chi_{j}\varphi^{,i}-{2\over 3}\delta^{i}_{j}\chi^{k}\varphi_{,k}\Big{)}\Big{]}\Big{\}}$
$\qquad-{c^{2}\over a^{2}(1+2\varphi)}\Big{[}{1\over 1+2\varphi}\Big{(}\nabla^{i}\nabla_{j}-{1\over 3}\delta^{i}_{j}\Delta\Big{)}\varphi+{1\over{\cal N}}\Big{(}\nabla^{i}\nabla_{j}-{1\over 3}\delta^{i}_{j}\Delta\Big{)}{\cal N}\Big{]}$
$\qquad=-{c^{2}\over a^{2}(1+2\varphi)^{2}}\Big{[}{3\over 1+2\varphi}\Big{(}\varphi^{,i}\varphi_{,j}-{1\over 3}\delta^{i}_{j}\varphi^{,k}\varphi_{,k}\Big{)}+{1\over{\cal N}}\Big{(}\varphi^{,i}{\cal N}_{,j}+\varphi_{,j}{\cal N}^{,i}-{2\over 3}\delta^{i}_{j}\varphi^{,k}{\cal N}_{,k}\Big{)}\Big{]}$
$\qquad+{c^{2}\over a^{4}{\cal N}^{2}(1+2\varphi)^{2}}\Big{[}{1\over 2}\Big{(}\chi^{i,k}\chi_{j,k}-\chi_{k,j}\chi^{k,i}\Big{)}+{1\over 1+2\varphi}\Big{(}\chi^{k,i}\chi_{k}\varphi_{,j}-\chi^{i,k}\chi_{j}\varphi_{,k}+\chi_{k,j}\chi^{k}\varphi^{,i}-\chi_{j,k}\chi^{i}\varphi^{,k}\Big{)}$
$\qquad+{2\over(1+2\varphi)^{2}}\Big{(}\chi^{i}\chi_{j}\varphi^{,k}\varphi_{,k}-\chi^{k}\chi_{k}\varphi^{,i}\varphi_{,j}\Big{)}\Big{]}+{8\pi G\over c^{2}}\overline{m}^{i}_{j},$ (103)
where
$\kappa\equiv 3{\dot{a}\over a}\Big{(}1-{1\over{\cal N}}\Big{)}-{1\over{\cal N}(1+2\varphi)}\Big{[}3\dot{\varphi}+{c\over a^{2}}\Big{(}\chi^{k}_{\;\;,k}+{\chi^{k}\varphi_{,k}\over 1+2\varphi}\Big{)}\Big{]},$ (104)
$\overline{K}^{i}_{j}\overline{K}^{j}_{i}={1\over a^{4}{\cal N}^{2}(1+2\varphi)^{2}}\Big{\{}{1\over 2}\chi^{i,j}\Big{(}\chi_{i,j}+\chi_{j,i}\Big{)}-{1\over 3}\chi^{i}_{\;\;,i}\chi^{j}_{\;\;,j}-{4\over 1+2\varphi}\Big{[}{1\over 2}\chi^{i}\varphi^{,j}\Big{(}\chi_{i,j}+\chi_{j,i}\Big{)}-{1\over 3}\chi^{i}_{\;\;,i}\chi^{j}\varphi_{,j}\Big{]}$
$\qquad+{2\over(1+2\varphi)^{2}}\Big{(}\chi^{i}\chi_{i}\varphi^{,j}\varphi_{,j}+{1\over 3}\chi^{i}\chi^{j}\varphi_{,i}\varphi_{,j}\Big{)}\Big{\}}.$ (105)
For fully nonlinear and exact formulation including the transverse (vector) and transverse-tracefree (tensor) modes to fully nonlinear and exact order, see Gong-2017 . The fluid quantities in Einstein’s equations are the collective ones, for example, $E=E_{m}+E_{\tau}$, $m_{i}=m_{i}^{m}+m_{i}^{\tau}$, $m_{ij}=m_{ij}^{m}+m_{ij}^{\tau}$, etc; for the $\tau$-fluid, fluid quantities are given in Eq. (99), and for ordinary fluid, scalar field, and Maxwell field, see Sec. IV.C in Hwang-Noh-EM-2023 .
The energy and momentum conservation equations give
${1\over a^{3}}\Big{[}a^{3}\Big{(}1+2\varphi\Big{)}^{3/2}E\Big{]}^{\displaystyle\cdot}+{1\over a}\Big{[}\Big{(}1+2\varphi\Big{)}^{1/2}\Big{(}{\cal N}c^{2}m^{i}+{c\over a}\chi^{i}E\Big{)}\Big{]}_{,i}=-{1\over a}\Big{(}1+2\varphi\Big{)}^{1/2}{\cal N}_{,i}c^{2}m^{i}$
$\qquad-\Big{(}1+2\varphi\Big{)}^{1/2}\Big{[}\Big{(}{\dot{a}\over a}+\dot{\varphi}+2{\dot{a}\over a}\varphi\Big{)}\delta^{ij}+{c\over a^{2}}\chi^{i,j}-{c\over a^{2}(1+2\varphi)}\Big{(}2\chi^{i}\varphi^{,j}-\delta^{ij}\chi^{k}\varphi_{,k}\Big{)}\Big{]}m_{ij},$ (106)
${1\over a^{4}}\Big{[}a^{4}\Big{(}1+2\varphi\Big{)}^{3/2}m_{i}\Big{]}^{\displaystyle\cdot}+{1\over a}\Big{[}\Big{(}1+2\varphi\Big{)}^{3/2}\Big{(}{\cal N}m_{i}^{j}+{c\chi^{j}m_{i}\over a(1+2\varphi)}\Big{)}\Big{]}_{,j}$
$\qquad={1\over a}\Big{(}1+2\varphi\Big{)}^{3/2}\Big{[}{{\cal N}\varphi_{,i}\over 1+2\varphi}S-{\cal N}_{,i}E-{c\over a}\Big{(}{\chi^{j}\over 1+2\varphi}\Big{)}_{,i}m_{j}\Big{]}.$ (107)
The conservation equations are valid for $\tau$-fluid and ordinary fluid, including scalar field as well as Maxwell field Hwang-Noh-EM-2023 . The equation of motion in Eq. (7) gives
${1\over c}{\partial\over\partial t}\Big{[}{a\over{\cal N}Q}\Big{(}{\cal N}\sqrt{1+2\varphi}{\cal J}_{,A}{Q^{,i}\over Q}\Big{)}_{,i}+{1\over 2}a^{3}(1+2\varphi)^{3/2}{\cal K}_{,Q}\Big{]}+{\partial\over\partial x^{i}}\Big{\{}\sqrt{1+2\varphi}{\cal J}_{,A}{Q^{,i}\over Q^{3}}\Big{(}{a\over c}\dot{Q}+{\chi^{j}Q_{,j}\over a(1+2\varphi)}\Big{)}$
$\qquad+{\chi^{i}\over a(1+2\varphi)}\Big{[}{1\over{\cal N}Q}\Big{(}{\cal N}\sqrt{1+2\varphi}{\cal J}_{,A}{Q^{,j}\over Q}\Big{)}_{,j}+{1\over 2}a^{2}(1+2\varphi)^{3/2}{\cal K}_{,Q}\Big{]}\Big{\}}=0.$ (108)
Using Eq. (99), we can show that, for $\tau$-fluid, Eq. (106) is the same as Eq. (108) and Eq. (107) is identically valid.
Ignoring the gravitational waves, Eqs. (100)-(108) are the complete set of exact perturbation equations. The conservation equations in (106)-(107) are valid for fluids as well as for $\tau$-field with the equation of state presented in Eq. (99); for the field we can also use the equation of motion in Eq. (108).
VI Discussion
A simple relativistic extension of MOND theory and its minimal modification to account for cosmological demands for DM was recently proposed by Blanchet, Marsat and Skordis Blanchet-Marsat-2011 ; Blanchet-Skordis-2024 ; we call it the BMS theory. The theory is summarized in Eqs. (1) and (2). Here, we derived equations of the PN approximation and perturbation theory in cosmological context.
The PN approximation of the BMS theory was presented in Flanagan-2023 ; Blanchet-Skordis-2024 in Minkowsky background, and the cosmological linear perturbation was presented in Blanchet-Skordis-2024 . Compared with previous works, here we presented equations valid to 1PN order in cosmology context, linear perturbations without imposing temporal gauge condition, and the fully-nonlinear and exact perturbation equations. We clarified the gauge issues in both formulations. We presented the 1PN approximation assuming $\tau=t$ as a physical condition and fully nonlinear and exact perturbations in $\tau=\tau(t)$ using a temporal gauge condition. General cases without imposing these conditions demand lengthy presentations, but the methods are the same.
In linear perturbations, we kept general ${\cal J}$ and ${\cal K}$ and derived Jeans criterion for a constraint on $\tau$-fluid. The Bekenstein and Milgrom formulation of MOND Bekenstein-Milgrom-1984 can be recovered by using ${\cal J}(A)$ in the 0PN limit of the PN approximation by taking $\sigma=0$ as a physical condition on $\tau$-field, thus $\tau=t$. The result is derived without assuming stationary condition mentioned in Flanagan-2023 ; Blanchet-Skordis-2024 .
Constraints on ${\cal K}(Q)$ follow from the background, perturbations, and the MOND requirements. (i) A demand for dust-like equation of state for $\tau$-field in the background constrains ${\cal K}$: see Eq. (20) for model-I. (ii) For model-II, a mass-like term appears for $n=1$, and the upper limit on this term gives a constraint in Eq. (48) which conflicts with the background constraint. These two constraints were presented in Blanchet-Skordis-2024 . (iii) A demand for $\tau$-fluid to behave as a pressureless fluid in linear density perturbation leads to a constraint. The Jeans scale is derived in Eq. (89) for general ${\cal J}$ and ${\cal K}$, and in Eq. (91) for model-I. Proper estimation of the effect of ${\cal K}$ in cosmological linear perturbations needs numerical study of the full system summarized below Sec. V.1.
In BMS theory, the two functions ${\cal J}$ and ${\cal K}$ are still free to accommodate the phenomena in MOND and cosmology. Future studies may design and constrain the functional forms using theories and observations. Other functional forms, especially ${\cal K}$ for cosmology were previously considered with successful fits to the power spectra of the large-scale matter and CMB Blanchet-Skordis-2024 ; Skordis-Zlosnik-2021 . The fully nonlinear and exact perturbation equations will be useful to study the leading nonlinear matter power spectrum as well as the non-Gaussian effects.
If successful, by adding two functions of a time-like vector field in a gradient form, the BMS theory encompasses the MOND phenomenon within Einstein’s gravity paradigm algebraically, although the local Lorentz invariance and the strong equivalence principle are broken conceptually. The latter differences may open new possibilities to test the theory.
Acknowledgments
H.N. was supported by the National Research Foundation (NRF) of Korea funded by the Korean Government (No.RS-2024-00333721 and No.2021R1F1A1045515). J.H. was supported by IBS under the project code, IBS-R018-D1.